OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult'a. The park offers half off all entry tickets during the off-peak season. The Jordan's paid a total of $69 for two adult tickets and one child's ticket to the waterpark during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket : 69 = x + 1/2 (x - 6) What is the regular price of a child's ticket?

OpenStudy (anonymous):

They give you the equation, what are you having trouble with?

OpenStudy (imstuck):

I'm guessing the problem is in solving for x right?

OpenStudy (anonymous):

69 = x + 1/2(x - 6) 69 = x + x/2 - 3

OpenStudy (anonymous):

Is that right IM?

OpenStudy (imstuck):

Yes, that's good...keep going... ; )

OpenStudy (anonymous):

69 = x + x/2 - 3 72 = x + x/2 2(72 = x + x/2) Correct?

OpenStudy (imstuck):

What do you get for x?

OpenStudy (anonymous):

144 = 2x + x 144 = 3x 48 = x

OpenStudy (imstuck):

If you solve it using your equation you will get it wrong. You need to find the common denominator between the x and the x/2.

OpenStudy (anonymous):

Okay so x / 1 + x / 2 2x / 2 + x / 2 2x + x

OpenStudy (anonymous):

...

OpenStudy (anonymous):

what is wrong with me?

OpenStudy (imstuck):

I get this:\[69=x+\frac{ 1 }{ 2 }x-3\]then\[72=\frac{ 2 }{ 2 }x+\frac{ 1 }{ 2 }x\]then\[72=\frac{ 3 }{ 2 }x\]

OpenStudy (anonymous):

That's because I said x/2 and not 1/2(x) because I derped the derp up.

OpenStudy (imstuck):

Multiply both sides by 2 to get rid of the fraction and it is now 144=3x and x = 48. Right. BUT...

OpenStudy (imstuck):

They want the price of a child's REGULAR admission ticket and x is the amount of an adult ticket at reduced price. Now we need to find the price of a regular child's ticket. Any ideas on that?

OpenStudy (imstuck):

They tell you in the problem that the special is half off the regular price when it is off-season. If $48 is the price during off-season, how do you find the price during peak season? What is your equation for that?

OpenStudy (anonymous):

You'd multiply it by 2.

OpenStudy (imstuck):

Yep. So now you have the regular price of an adult ticket. What do they tell you about the price of a child's ticket in regards to the adult's ticket? How much less is the child's ticket?

OpenStudy (anonymous):

if x is equal to 48 and a regular child's ticket it 6$ less than an adults ticket wouldn't you just subtract 6 and get 42?

OpenStudy (anonymous):

The answer choices are $38 $42 $48 and $66

OpenStudy (mathstudent55):

\(69 = x + 1/2 (x - 6)\) Multiply both sides by 2: 138 = 2x + x - 6 138 = 3x - 6 144 = 3x x = 48

OpenStudy (anonymous):

48 was the wrong answer

OpenStudy (mathstudent55):

48 is x. I didn't say the answer is x. x is the price for an adult. The price for a child is $48 - $6 = $42. I was just showing how I solve the equation because I wanted to show the way I get rid of the denominator early on.